# The Coase Theorem and Ideal Exchanges

### v2

@@ 1. Introduction
Prior to Ronald Coase’s work, The Problem of Social Cost, the conventional view—which Coase himself attributes to Arthur Pigou [1, p. 1]—held that externalities could be corrected through state intervention: namely, through taxes or subsidies that are designed to deter people from either over-consuming goods associated with negative externalities or under-consuming goods associated with positive externalities [2]. Coase reoriented the debate from one that relied solely on governmental regulation to one that recognized the potential efficiency of private negotiation in resolving externalities. In particular, Coase reasoned that externalities resulted from ill-defined property rights; provided that property rights are well-defined and each party could negotiate with the other without hindrance, resources would be distributed away from those who value them less and towards those who value them more irrespective of how the rights are initially distributed [1, p. 8, para. 2]. This proposition, although informal, was characterized as a theorem in George Stigler’s Theory of Price:
@@ 2. Preliminaries
Every Pareto optimal distribution occurs when the indifference curves for both agents are tangent to each other. We can see this by considering the intersection at $\omega_{A}$ or, equivalently, $\omega_{B}$. Since $A$ has an excess of good $x$ and $B$ has an excess of good $y$, it would be possible for $A$ to give up just enough of $x$ and receive just enough of $y$ so that she is indifferent between $\omega_{A}$ and $\Omega$, while $B$’s indifference curve progresses to a point where it is no longer possible to gain any further utility without damaging $A$’s material interests; the opposite is true at $\omega_{A}^{\prime}$. If we assume that exchanges can occur iff they are mutually beneficial, then the negotiated outcome is strictly inside the lens formed by the two indifference curves. Through continual negotiations, the size of the lens shrinks until it is no longer possible to make a mutually beneficial trade. [^4] This is a standard description of how negotiations occur. It is notably present in Buchanan and Tullock’s foundational work on public choice theory, The Calculus of Consent: Logical Foundations of Constitutional Democracy [8, p. 100]. Hence, the ultimate outcome, although different from $\Omega$, would still be Pareto optimal. A stronger variant of the theorem in narrow cases where a medium of exchange is involved was proven by Leonid Hurwicz, who demonstrated that the invariance of the final distribution of a particular resource necessarily depended on the assumption of quasi-linear preferences (i.e., $\mathcal{U}_{A}(Q_{x},m)=v(Q_{x})+m$, where $m$ is a “numeraire” good) [9].
It is a striking fact that the preceding methods are standard in microeconomic theory yet are obviously mathematically unnatural for the purposes of modeling economic behavior. Francis Ysidro Edgeworth, after whom the Edgeworth box is named, introduced the concept of indifference curves in Mathematical Psychics (MP), a foundational work in contemporary microeconomic theory [10, p. 28]. The broad purpose of MP was to apply mathematical techniques to the moral sciences; in doing so, Edgeworth introduced utility measurements by assigning real numbers to represent individuals’ levels of satisfaction or happiness, facilitating the comparison and aggregation of utility across different people. Furthermore, MP readily and implicitly utilizes the concept of a continuum—suggesting infinitely divisible goods or utility levels—without being concerned with foundational aspects of real analysis. On a historical note, Dedekind’s Stetigkeit und irrationale Zahlen was published in 1872, merely 9 years before MP [11]; it is therefore likely that Edgeworth was unaware of these developments in the philosophical foundations of mathematics and their effects on the mathematical foundations of economics.
@@
The attempts to formalize mathematics and make precise what the axioms are were never thought of as attempts to explain the rules of logic, but rather, to write down these rules and axioms which appeared to correspond to what contemporary mathematicians were using. [12, p. 1074]
@@ 3. Foundations
Let $\mathbb{A}$ be a set of agents of cardinality $n\in\mathbb{N}$ and $R$ be a set of resources of cardinality $m\in\mathbb{N}$. The distribution of resources across each instance of trade $t\in\mathbb{N}$ can be characterized as a mapping $\mathcal{I}_{t}:\mathbb{A}\to\mathcal{P}(R)$ such that the following two conditions are satisfied:
@@ 4. Secondary Results
$\forall\,\big{(}\mathcal{I}_{0}(a_{i})\big{)}_{i=0}^{n-1}\in\mathcal{P}(R)^{n}\,\exists\,t\in\mathbb{N}\,\forall\,k\in\mathbb{N}\,\exists\,\big{(}\alpha_{i}\big{)}_{i=0}^{n-1}\in\mathcal{P}(R)^{n}\Big{[}\big{(}\mathcal{I}_{t+k}(a_{i})\big{)}_{i=0}^{n-1}=\big{(}\alpha_{i}\big{)}_{i=0}^{n-1}\Big{]}.$
@@ 6. Philosophical & Concluding Remarks
We now turn to the primary disproof in 4.6 where we offer an explicit counterexample to Proposition 4.5. Evidently, it would seem that the exchange mechanism prescribed by (C1) is too strong, and it has been argued that the inflexibility of a double coincidence of wants is a problem in markets where exchanges are conducted through bartering. The problem is certainly not unique to such societies and has not fully disappeared in markets where there exists a dominant medium of exchange (i.e., where it has only morphed into the need for a double coincidence of value.) To see this, we may turn to 4.6 and reinterpret $z$ as a $20\$$ bill, whereas $x$ and $y$ continue to stand for material goods. $a_{0}$ still prefers $y$ over $x$ and $x$ over $20\$$: $a_{1}$ prefers $x$ over $20\$$ and $20\$$ over $y$, while $a_{2}$ prefers $20\$$ over $y$ and $y$ over $x$. If the initial allocation were ($\big{(}\{x\},\{20\$\},\{y\}\big{)}$, then $a_{1}$ cannot purchase $x$ from $a_{0}$, who values it over $20\$$. It follows from 4.6 that no exchange occurs and the distribution is suboptimal. This is a mathematically trivial demonstration, yet its philosophical significance for economic theory lies in the observation that a market exchange should not be viewed as a private activity among only two parties. If it were, then it would be irrational for $a_{1}$ to purchase $y$ from $a_{2}$ with $20\$$, as that would make her worse off immediately. But if she chooses to do so and subsequently barters $y$ for $x$ with $a_{0}$, then everyone is made better off. Naturally, it would be more efficient if all three agents recognized this, and immediately arrived at a better distribution through a 3-party agreement. The presence of a common medium of exchange has a non-trivial effect if we assume that preferences are incomplete over resources but complete over the medium of exchange; we reject this two-sorted-ness as it is irrational and would contradict (B3). Furthermore, it is inconsistent with (B2) because the defining quality of a medium of exchange is its ability to measure the value of an item. If we can compare an item to a quantity of the medium, and compare that quantity with the quantity we derive from some other comparable item, then we should be able to compare the items themselves as though we lived in a bartering economy. [^7] Let $R$ be a set of resources such that it consists of a unique multiset of identical goods $\mathcal{N}\subseteq R$ for which every agent has quasi-linear preferences: $\forall a_{q}\in\mathbb{A}\forall A,B\in\mathcal{P}(R)\big{[}(((A\setminus(A\cap B)\subseteq\mathcal{N})\land(B\setminus(A\cap B)\subseteq\mathcal{N}))\land(|A\setminus(A\cap B)|<|B\setminus(A\cap B)|))\implies(A,B)\in W(a_{q})\big{]}.$ The existence of this numeraire does not solve the predicament since agents may freely elect to not have a price tag on particular material resources. In other words, no amount of money would be enough to compel them to sell off what they own (i.e., $\exists a_{q}\in\mathbb{A}\exists A\in\mathcal{P}(R)\forall B\subseteq\mathcal{N}[(B,A)\in W(a_{q})]$). If one were to contend that a price tag can be assigned to every bundle of resources in $\mathcal{P}(R)$ (i.e., $\forall a_{q}\in\mathbb{A}\,\forall A\in\mathcal{P}(R)\,\exists B\subseteq\mathcal{N}[(A,B)\in W(a_{q})]$) then this would call into question the value of the medium. Evidently, the numeraire’s worth is contingent on the value of the goods it can buy: it cannot be pursued as a bundle of resources in its own right, rather, there must exist some bundle of resources such that that bundle is strictly preferred over the numeraire. But this would create a vicious cycle: since there must exist a price tag for that bundle, there must also be a different corresponding bundle for that price.

### v4

@@ 1. Introduction
Prior to Ronald Coase’s work, The Problem of Social Cost, the conventional view—which Coase himself attributes to Arthur Pigou [4, p. 1]—held that externalities could be corrected through state intervention: namely, through taxes or subsidies that are designed to deter people from either over-consuming goods associated with negative externalities or under-consuming goods associated with positive externalities [11]. Coase reoriented the debate from one that relied solely on governmental regulation to one that recognized the potential efficiency of private negotiation in resolving externalities. In particular, Coase reasoned that externalities resulted from ill-defined property rights; provided that property rights are well-defined and each party could negotiate with the other without hindrance, resources would be distributed away from those who value them less and towards those who value them more irrespective of how the rights are initially distributed [4, p. 8, para. 2]. This proposition, although informal, was characterized as a theorem in George Stigler’s Theory of Price:
@@ 2. Preliminaries
Every Pareto optimal distribution occurs when the indifference curves for both agents are tangent to each other. We can see this by considering the intersection at $\omega_{A}$ or, equivalently, $\omega_{B}$. Since $A$ has an excess of good $x$ and $B$ has an excess of good $y$, it would be possible for $A$ to give up just enough of $x$ and receive just enough of $y$ so that she is indifferent between $\omega_{A}$ and $\Omega$, while $B$’s indifference curve progresses to a point where it is no longer possible to gain any further utility without damaging $A$’s material interests; the opposite is true at $\omega_{A}^{\prime}$. If we assume that exchanges can occur iff they are mutually beneficial, then the negotiated outcome is strictly inside the lens formed by the two indifference curves. Through continual negotiations, the size of the lens shrinks until it is no longer possible to make a mutually beneficial trade. [^4] This is a standard description of how negotiations occur. It is notably present in Buchanan and Tullock’s foundational work on public choice theory, The Calculus of Consent: Logical Foundations of Constitutional Democracy [2, p. 100]. Hence, the ultimate outcome, although different from $\Omega$, would still be Pareto optimal. A stronger variant of the theorem in narrow cases where a medium of exchange is involved was proven by Leonid Hurwicz, who demonstrated that the invariance of the final distribution of a particular resource necessarily depended on the assumption of quasi-linear preferences (i.e., $\mathcal{U}_{A}(Q_{x},m)=v(Q_{x})+m$, where $m$ is a “numeraire” good) [9].
It is a striking fact that the preceding methods are standard in microeconomic theory yet are obviously mathematically unnatural for the purposes of modeling economic behavior. Francis Ysidro Edgeworth, after whom the Edgeworth box is named, introduced the concept of indifference curves in Mathematical Psychics (MP), a foundational work in contemporary microeconomic theory [8, p. 28]. The broad purpose of MP was to apply mathematical techniques to the moral sciences; in doing so, Edgeworth introduced utility measurements by assigning real numbers to represent individuals’ levels of satisfaction or happiness, facilitating the comparison and aggregation of utility across different people. Furthermore, MP readily and implicitly utilizes the concept of a continuum—suggesting infinitely divisible goods or utility levels—without being concerned with foundational aspects of real analysis. On a historical note, Dedekind’s Stetigkeit und irrationale Zahlen was published in 1872, merely 9 years before MP [6]; it is therefore likely that Edgeworth was unaware of these developments in the philosophical foundations of mathematics and their effects on the mathematical foundations of economics.
@@
The attempts to formalize mathematics and make precise what the axioms are were never thought of as attempts to explain the rules of logic, but rather, to write down these rules and axioms which appeared to correspond to what contemporary mathematicians were using. [3, p. 1074]
@@ 3. Foundations
Let $\mathbb{A},R\neq\varnothing$ with $|\mathbb{A}|=n,\;|R|=m$ and $n,m\in\mathbb{N}$. The distribution of resources across each instance of trade $t\in\mathbb{N}$ can be characterized as a mapping $\mathcal{I}_{t}:\mathbb{A}\to\mathcal{P}(R)$ such that the following two conditions are satisfied:
@@ 4. Secondary Results
Due to formal necessity, throughout the following we shall work within a fixed economy $\mathcal{E}$ on which every $\mathcal{I}$ operates, where $\mathcal{E}$ is defined as the 3-tuple $(\mathbb{A},R,W)$.
@@
$\forall\,\big{(}\mathcal{I}_{0}(a_{i})\big{)}_{i=0}^{n-1}\neq\big{(}\mathcal{I}^{\prime}_{0}(a_{i})\big{)}_{i=0}^{n-1}\in\mathcal{P}(R)^{n}\,\exists\,\big{(}\alpha_{i}\big{)}_{i=0}^{n-1}\in\mathcal{P}(R)^{n}\,$
$\quad\exists t,t^{\prime}\in\mathbb{N}\;\;\forall k,k^{\prime}\in\mathbb{N}\,\Big{[}\big{(}\mathcal{I}_{t+k}(a_{i})\big{)}_{i=0}^{n-1}=\big{(}\alpha_{i}\big{)}_{i=0}^{n-1}=\big{(}\mathcal{I}_{t^{\prime}+k^{\prime}}(a_{i})\big{)}_{i=0}^{n-1}\big{]}.$
@@ 6. Philosophical & Concluding Remarks
We now turn to the primary disproof in 4.6 where we offer an explicit counterexample to Proposition 4.5. Evidently, it would seem that the exchange mechanism prescribed by (C1) is too strong, and it has been argued that the inflexibility of a double coincidence of wants is a problem in markets where exchanges are conducted through bartering. The problem is certainly not unique to such societies and has not fully disappeared in markets where there exists a dominant medium of exchange (i.e., where it has only morphed into the need for a double coincidence of value.) To see this, we may turn to 4.6 and reinterpret $z$ as a $20\$$ bill, whereas $x$ and $y$ continue to stand for material goods. $a_{0}$ still prefers $y$ over $x$ and $x$ over $20\$$: $a_{1}$ prefers $x$ over $20\$$ and $20\$$ over $y$, while $a_{2}$ prefers $20\$$ over $y$ and $y$ over $x$. If the initial allocation were ($\big{(}\{x\},\{20\$\},\{y\}\big{)}$, then $a_{1}$ cannot purchase $x$ from $a_{0}$, who values it over $20\$$. It follows from 4.6 that no exchange occurs and the distribution is suboptimal. This is a mathematically trivial demonstration, yet its philosophical significance for economic theory lies in the observation that a market exchange should not be viewed as a private activity among only two parties. If it were, then it would be irrational for $a_{1}$ to purchase $y$ from $a_{2}$ with $20\$$, as that would make her worse off immediately. But if she chooses to do so and subsequently barters $y$ for $x$ with $a_{0}$, then everyone is made better off. Naturally, it would be more efficient if all three agents recognized this, and immediately arrived at a better distribution through a 3-party agreement. The presence of a common medium of exchange has a non-trivial effect if we assume that preferences are incomplete over resources but complete over the medium of exchange; we reject this two-sorted-ness as it is irrational and would contradict (B3). Furthermore, it is inconsistent with (B2) because the defining quality of a medium of exchange is its ability to measure the value of an item. If we can compare an item to a quantity of the medium, and compare that quantity with the quantity we derive from some other comparable item, then we should be able to compare the items themselves as though we lived in a bartering economy. [^7] Let $R$ be a set of resources such that it consists of a unique multiset of identical goods $\mathcal{N}\subseteq R$ for which every agent has quasi-linear preferences: $\forall a_{q}\in\mathbb{A}\forall A,B\in\mathcal{P}(R)\big{[}(((A\setminus(A\cap B)\subseteq\mathcal{N})\land(B\setminus(A\cap B)\subseteq\mathcal{N}))\land(|A\setminus(A\cap B)|<|B\setminus(A\cap B)|))\implies(A,B)\in W(a_{q})\big{]}.$ The existence of this numeraire does not solve the predicament since agents may freely elect to not have a price tag on particular material resources. In other words, no amount of money would be enough to compel them to sell off what they own (i.e., $\Diamond\,\exists a_{q}\in\mathbb{A}\exists A\in\mathcal{P}(R)\forall B\subseteq\mathcal{N}[(B,A)\in W(a_{q})]$). If one were to contend that a price tag can be assigned to every bundle of resources in $\mathcal{P}(R)$ (i.e., $\Box\,\forall a_{q}\in\mathbb{A}\,\forall A\in\mathcal{P}(R)\,\exists B\subseteq\mathcal{N}[(A,B)\in W(a_{q})]$) then this would call into question the value of the medium. Evidently, the numeraire’s worth is contingent on the value of the goods it can buy: it cannot be pursued as a bundle of resources in its own right, rather, there must exist some bundle of resources such that that bundle is strictly preferred over the numeraire. But this would create a vicious cycle: since there must exist a price tag for that bundle, there must also be a different corresponding bundle for that price.